# Patterns Over Principles: The Fragility of Inductive Reasoning in LLMs under Noisy Observations

### v1

@@ 6 Related Work
The study of inductive reasoning has been a long-standing focus across multiple disciplines. Early work Heit (2000) established foundational properties of inductive reasoning. In cognitive science, induction is considered a process of probabilistic belief updating within the Bayesian paradigm Tenenbaum et al. (2011). Human cognition and learning are explained through the integration of prior knowledge with observed data to compute posterior distributions. Comparative studies Lake et al. (2015, 2017) have highlighted the contrast between human learners and machine intelligence. With the advent of pre-trained language models Brown et al. (2020), research on inductive reasoning has shifted from domain-specific and neural formulation Tian et al. (2020); Odena et al. (2021); Sablé-Meyer et al. (2022) to natural language. Initial approaches Alet et al. (2021); Gendron et al. (2024); Mirchandani et al. (2023); Yang et al. (2024) predominantly relied on input-output (IO) prompting, which evaluates model performance on unseen examples without explicit rule articulation. However, this paradigm overlooks the internal rule-inference process, as it conflates rule induction with rule execution capabilities. Recent efforts Wang et al. (2024) align more closely with our work by explicitly generating intermediate rules. Qiu et al. (2024) proposed a thorough evaluation containing noisy conditions, whose findings partially overlap with ours, but their analysis is confined to a single dataset and prioritized accuracy metrics over consistency. We further explore the challenges posed by counterfactual tasks and uncover the limitations of LLMs in inductive reasoning.
@@
In this work, robustness refers to the ability to maintain consistent performance under imperfections or counterfactual scenarios Elazar et al. (2021), which is intrinsically linked to the diversity and unpredictability of generation process Zhang et al. (2023); Huang et al. (2025b). Prior works into LLM behavioral variability exhibit inconsistent performance across time Tu et al. (2024); Chen et al. (2023a). Recent works focus on understanding their internal reasoning mechanisms by altering conditions, such as changing numerical values in mathematical tasks Mirzadeh et al. (2024); Huang et al. (2025a) or examining performance on counterfactual tasks Wu et al. (2024). Zhou et al. (2024) investigates the impact of noisy rationales on model performance. These studies primarily measure robustness through global accuracy changes, overlooking inconsistencies at the instance level. Our work addresses this gap by introducing a consistency score that quantifies intra-task stability, providing a more granular view of model robustness.
@@ 7 Conclusion
We discuss the limitations of our work here: Our evaluation focuses on highly formalized and symbolic tasks. Real-world inductive reasoning often involves ambiguous rules like social norms from text or visual patterns, which are not captured by our formalism. The tasks in Arithmetic and Cryptography may be relatively simple and lack diversity. Expanding task diversity could reveal deeper limitations in LLMs’ capabilities. A large-scale assessment of human performance across all tasks under varying levels of noise is beyond the scope of this study, which poses certain limitations in comparing human reasoning patterns. The purpose of our study is to explore the robustness of LLMs’ inductive reasoning capabilities, so we do not specifically tune the hyperparameters and prompt templates.
@@
The datasets we used are all publicly available, and our research does not involve any personal information. All data is generated by programs. Therefore, we anticipate that this paper does not raise any ethical concerns. We use ChatGPT to paraphrase some sentences.
@@ Appendix A Details on Evaluation Pipeline
In this section, we provide detailed information on the evaluation pipeline, including the data construction and the performance assessment. For Arithmetic, we generate the base-7, base-8, and base-9 tasks by randomly sampling two two-digit numbers in the corresponding base, and then we check whether there is a carry-over in the addition process. If there is no carry-over, we regenerate the numbers. The noisy examples are generated in base-10. For Cryptography, we randomly select words of appropriate lengths from the NLTK Word Lists corpus111https://www.nltk.org/nltk_data/, and then we encrypt the words using the Caesar, Atbash, and Keyboard ciphers. The noisy examples are generated by randomly replacing the letters in the output with other letters. For List Functions, we first write the corresponding rule functions for each task, and then we automatically generate the input data with appropriate lengths and ranges. The inputs are generated by randomly sampling numbers from a specific range with some constraints. . During the data synthesis process, we attach manual supervision to ensure that the generated data can correctly induce the rules. During the evaluation process, we use the inputs in the test set as the input for rule execution. We evaluate the model’s performance using exact match. If the model’s output is correct on all the test set examples, we consider the model to have successfully induced the rule. If the model fails to output a valid programmatic rule or the program contains an infinite loop or errors, we consider it a failure.
@@ B.1 Experimental Settings
For robustness under different noise levels, we run each experiment three times and report the mean and standard deviation of the results to avoid randomness. Except for the self-consistency (SC) and sample-steered rule refinement (SRR) that require diverse generations, we set the temperature to for all models to ensure reproducibility. For SC and SRR, the temperature is set to , consistent with the original work of SC Wang et al. (2023). The positions of noise in seen examples are random to avoid positional bias Lu et al. (2022). In the implementation, we choose subsets for SRR by splitting the seen examples into two parts. The number of iterations is set to for both SR and SRR.
@@
All experiments are conducted through the official OpenAI222https://openai.com/api/ and DeepSeek333https://platform.deepseek.com/usage API platform. For the GPT-4o-mini and GPT-4o models, we spend about USD in total, and for the DeepSeek-V3 and DeepSeek-R1 models, we spend about USD in total.
B.2 Prompts and Failure Cases
For the Direct Output setting, we restrict the model to output the rule directly without any additional output, as shown in Table 7. For the chain-of-thought and self-consistency setting, we use the instruction in Kojima et al. (2022). For the self-refine and sample-steered rule refinement setting, we use the chain-of-thought prompt as the initial prompt; the iterative prompts are shown in Table 9 and Table 10, respectively.
We provide some representative failure cases in the evaluation. For the Arithmetic task, the model fails to solve the base-7 and base-9 tasks and misinterprets the rule as the base-8 addition or the decimal sum with a constant. For the Cryptography task, the model fails to solve the Atbash and tries to explain it as shifts. The responses of the models are shown in Table 11.
Prompt for Direct Output |
Please generate a rule that maps the following inputs to their corresponding outputs using a Python function. The input is a list of integers. The output is also a list of integers. Note that some examples may be wrong, and you should take this into account when proposing the rule. |
{examples} |
Please format your Python function as follows: |
“‘python |
def fn(x): |
# Your code here |
Your response should only include the function definition, not the function call or any other information. |
Prompt for Chain-of-Thought |
Please generate a rule that maps the following inputs to their corresponding outputs using a Python function. The input is a list of integers. The output is also a list of integers. Note that some examples may be wrong, and you should take this into account when proposing the rule. |
{examples} |
Please format your Python function as follows: |
“‘python |
def fn(x): |
# Your code here |
Think step-by-step and explain your reasoning. Your response should include your thought process and the function definition without the function call. |
Prompt for Feedback Generation in Self-Refine |
You have generated a rule that maps the following inputs to their corresponding outputs using a Python function. The input is a list of integers. The output is also a list of integers. |
{examples} |
In the last step, your rule is: |
“‘python |
{rule} |
Give some feedback on the rule you have generated, like how can it be improved, what is wrong with it, etc. |
Your response should only include the feedback. If you think the rule is good enough, your response should be“NO FEEDBACK” without other information. Note that some examples may be wrong, and you should take this into account when proposing the feedback. |
Iteration |
You have generated a rule that maps the following inputs to their corresponding outputs using a Python function. The input is a list of integers. The output is also a list of integers. Note that some examples may be wrong, and you should take this into account when proposing the rule. |
{examples} |
In the last step, your rule is: |
“‘python |
{rule} |
The feedback you have given is: |
{feedback} |
Generate a new rule that maps the given inputs to their corresponding outputs using a Python function. Please format your rule as follows: |
“‘python |
def fn(x): |
# Your code here |
Think step-by-step and explain your reasoning. Your response should include your thought process and the function definition without the function call. |
Prompt for Iteration in SRR |
You have generated a rule that maps the following inputs to their corresponding outputs using a Python function. The input is a list of integers. The output is also a list of integers. Note that some examples may be noisy, and you should take this into account when proposing the rule. In the last step, your rule is |
“‘python |
{rule} |
But this rule is not correct. It works for the following examples: |
{right_examples} |
However, it does not work for the following examples: |
{wrong_examples} Generate a new rule that maps the given inputs to their corresponding outputs using a Python function. Please format your rule as follows: |
“‘python |
def fn(x): |
# Your code here |
Think step-by-step and explain your reasoning. Your response should include your thought process and the function definition without the function call. You can either modify the existing rule or propose a new one. |
Failure Cases for Arithmetic |
Octonary Confusion Error (in base-7): |
To solve this problem, we need to identify a pattern or rule that maps the given input expressions to their corresponding outputs. Let’s analyze the examples provided: First, let’s verify if the outputs are simply the sum of the inputs in base 10:- For ‘34 + 62‘, the sum is ‘96‘, but the output is ‘126‘.- For ‘56 + 30‘, the sum is ‘86‘, but the output is ‘116‘.- For ‘22 + 25‘, the sum is ‘47‘, but the output is ‘50‘… Clearly, the outputs are not the direct sums of the inputs in base 10. Let’s consider the possibility that the numbers are in base 8 (octal):- Convert the inputs to base 10, perform the addition, and then convert the result back to base 8. Let’s verify this hypothesis: 1. 34 + 62 in base 8: - ‘34‘ in base 8 is ‘3*8 + 4 = 28‘ in base 10. - ‘62‘ in base 8 is ‘6*8 + 2 = 50‘ in base 10. - Sum in base 10: ‘28 + 50 = 78‘. - Convert ‘78‘ to base 8: ‘78 // 8 = 9‘ remainder ‘6‘, so ‘78‘ in base 8 is ‘116‘.2. 56 + 30 in base 8: - ‘56‘ in base 8 is ‘5*8 + 6 = 46‘ in base 10. - ‘30‘ in base 8 is ‘3*8 + 0 = 24‘ in base 10. - Sum in base 10: ‘46 + 24 = 70‘. - Convert ‘70‘ to base 8: ‘70 // 8 = 8‘ remainder ‘6‘, so ‘70‘ in base 8 is ‘106‘. 3. 22 + 25 in base 8: - ‘22‘ in base 8 is ‘2*8 + 2 = 18‘ in base 10. - ‘25‘ in base 8 is ‘2*8 + 5 = 21‘ in base 10. - Sum in base 10: ‘18 + 21 = 39‘. - Convert ‘39‘ to base 8: ‘39 // 8 = 4‘ remainder ‘7‘, so ‘39‘ in base 8 is ‘47‘. Following this pattern, we can see that the outputs are indeed the sums of the inputs when interpreted as base 8 numbers and then converted back to base 8. Let’s implement this in a Python function… |
Decimal Confusion Error (in base-9): |
To solve this problem, we need to identify a consistent rule or pattern that maps the given input expressions to their corresponding outputs. Let’s analyze the provided examples: 1. : The correct sum is 136, but the output is 147. 2. : The correct sum is 122, but the output is 132. 3. : The correct sum is 93, but the output is 104. 4. : The correct sum is 52, but the output is 53. 5. : The correct sum is 127, but the output is 137. 6. : The correct sum is 103, but the output is 114. 7. : The correct sum is 128, but the output is 138. 8. : The correct sum is 80, but the output is 81. 9. : The correct sum is 103, but the output is 114. 10. : The correct sum is 107, but the output is 117. From these examples, it appears that the output is consistently 10 more than the correct sum of the two numbers. This suggests that the rule is to calculate the sum of the two numbers and then add 10 to the result. Let’s define a Python function to implement this rule… |
Failure Case for Cryptography |
Shifts Error: |
To solve this problem, we need to identify a consistent rule or transformation that maps each input string to its corresponding output string. Let’s analyze the given examples to find a pattern: 1. cuissard -> xfrhhziw 2. tripsis -> girkhrh 3. Maccabaeus -> Nzxxzyzvfh 4. jocundness -> qlxfmwmvhh 5. chiro -> xsril 6. coapt -> xlzkg 7. Phecda -> Ksvxwz 8. unifilar -> fmrurozi 9. Cordaites -> Xliwzrgvh 10. Wendell -> Dvmwvoo ### Observations: 1. Character Shifts : The transformation seems to involve shifting characters. For example, ’c’ in "cuissard" becomes ’x’ in "xfrhhziw". This suggests a possible Caesar cipher or similar character shift. 2. Case Sensitivity : The case of the letters is preserved. For instance, ’M’ in "Maccabaeus" becomes ’N’ in "Nzxxzyzvfh", and ’P’ in "Phecda" becomes ’K’ in "Ksvxwz". 3. Consistent Shifts : By examining the shifts: - ’c’ to ’x’ is a shift of 23 positions backward in the alphabet. - ’t’ to ’g’ is a shift of 13 positions backward. - ’M’ to ’N’ is a shift of 1 position forward. - ’j’ to ’q’ is a shift of 7 positions forward. - ’P’ to ’K’ is a shift of 21 positions backward. - ’u’ to ’f’ is a shift of 11 positions backward. - ’C’ to ’X’ is a shift of 21 positions backward. - ’W’ to ’D’ is a shift of 21 positions backward. 4. Pattern Recognition : The shifts are not consistent across all examples, suggesting that some examples might be incorrect or that multiple rules are being applied. However, a common shift that appears frequently is 21 positions backward (e.g., ’c’ to ’x’, ’P’ to ’K’, ’C’ to ’X’, ’W’ to ’D’). ### Hypothesis: Given the frequent appearance of a 21-position backward shift, let’s hypothesize that the rule involves shifting each character 21 positions backward in the alphabet. This would mean: - For lowercase letters, ’a’ becomes ’v’, ’b’ becomes ’w’, …, ’z’ becomes ’u’. - For uppercase letters, ’A’ becomes ’V’, ’B’ becomes ’W’, …, ’Z’ becomes ’U’. ### Implementation: Let’s implement this hypothesis in a Python function… |

### v2

@@ 6 Related Work
Early work Heit (2000) established foundational properties of inductive reasoning. In cognitive science, induction is considered a process of probabilistic belief updating within the Bayesian paradigm Tenenbaum et al. (2011). Human cognition and learning are explained through the integration of prior knowledge with observed data to compute posterior distributions. Comparative studies Lake et al. (2015, 2017) have highlighted the contrast between human learners and machine intelligence. With the advent of pre-trained language models Brown et al. (2020), research on inductive reasoning has shifted from domain-specific and neural formulation Tian et al. (2020); Odena et al. (2021); Sablé-Meyer et al. (2022) to natural language. Initial approaches Alet et al. (2021); Gendron et al. (2024); Mirchandani et al. (2023); Yang et al. (2024) predominantly relied on input-output (IO) prompting, which evaluates model performance on unseen examples without explicit rule articulation. However, this paradigm overlooks the rule-inference process, as it conflates rule induction with rule execution capabilities. Recent efforts Wang et al. (2024); Qiu et al. (2024) explicitly evaluate intermediate rules, and our settings and methods are similar to theirs, despite the sampling part in initialization and iteration. Specifically, Qiu et al. (2024) also proposed an evaluation on noisy conditions, but their analysis is confined to a single dataset and prioritized accuracy metrics. We further explore the consistency and draw a different conclusion that LLMs’ inductive reasoning capabilities are not robust and generalizable.
@@
In this work, robustness refers to the ability to maintain consistent performance under imperfections or counterfactual scenarios Elazar et al. (2021), which is intrinsically linked to the diversity and unpredictability of the generation process Zhang et al. (2023); Huang et al. (2025b). Prior works on LLM behavioral variability exhibit inconsistent performance across time Tu et al. (2024); Chen et al. (2023a). Recent works focus on understanding their internal reasoning mechanisms by altering conditions, such as changing numerical values in mathematical tasks Mirzadeh et al. (2024); Huang et al. (2025a) or examining performance on counterfactual tasks Wu et al. (2024). Zhou et al. (2024) investigates the impact of noisy rationales on model performance. These studies primarily measure robustness through global accuracy changes, overlooking inconsistencies at the instance level. Our work addresses this gap by introducing a consistency score that quantifies intra-task stability, providing a more granular view of model robustness.
@@ 7 Conclusion
We discuss the limitations of our work as follows. First, the general goal of our research is to investigate the relationship between a model’s reasoning and its understanding, specifically examining whether correctly solving a problem indicates true comprehension and reasoning of the underlying issue. To this end, our evaluation primarily focuses on highly formalized and symbolic tasks that are easily verifiable. However, real-world inductive reasoning often involves ambiguous rules like social norms from text or visual patterns, which are not captured by our formalism. Second, the tasks in Arithmetic and Cryptography may be relatively simple and unified, potentially limiting insights into the capabilities of large language models. Expanding task diversity could uncover deeper limitations in LLMs’ reasoning abilities. Finally, conducting a large-scale assessment of human performance across all tasks under varying levels of noise was beyond the scope of this study, which restricts our ability to compare human and model reasoning patterns comprehensively.
@@
We discuss the ethical considerations of our work in the context of the following aspects: (1) Data collection and use. We use publicly available datasets and self-generated data for evaluation. We ensure that the data is only used for academic research purposes and no personal data is involved. The List Functions dataset is released with Apache-2.0 license111https://www.apache.org/licenses/LICENSE-2.0 in Big-bench bench authors (2023). We strictly follow the license terms and conditions. (2) LLMs API. We comply with the terms of service of the LLMs API providers strictly, maintaining fair use. (3) Transparency and reproducibility. We provide detailed descriptions of our methods, datasets, and evaluation metrics to ensure transparency. Our code and evaluation results are publicly available. (4) AI assistance. We use ChatGPT to paraphrase some sentences.
Acknowledgement
We thank all the anonymous reviewers and meta reviewers for their valuable comments. The authors of this paper were supported by the ITSP Platform Research Project (ITS/189/23FP) from ITC of Hong Kong, SAR, China, and the AoE (AoE/E-601/24-N), the RIF (R6021-20) and the GRF (16205322) from RGC of Hong Kong, SAR, China.
@@ Appendix A Details on Evaluation Pipeline
In this section, we provide detailed information on the evaluation pipeline, including the data construction and the performance assessment.
For Arithmetic, we generate the base-7, base-8, and base-9 tasks by randomly sampling two two-digit numbers in the corresponding base, and then we check whether there is a carry-over in the addition process. If there is no carry-over, we regenerate the numbers. The noisy examples are generated in base-10. This choice reflects a common human-like calculation error—defaulting to the familiar decimal system—making it a “meaningful” and realistic form of noise in a mathematical context.
For Cryptography, we randomly select words of appropriate lengths from the NLTK Word Lists corpus222https://www.nltk.org/nltk_data/, and then we encrypt the words using the Caesar, Atbash, and Keyboard ciphers. The noisy examples are generated by randomly replacing the letters in the output with other letters. This mimics common errors in text processing or encryption, such as typographical mistakes or transmission noise, which are naturally occurring in symbolic domains.
For List Functions, we first write the corresponding rule functions for each task, and then we automatically generate the input data with appropriate lengths and ranges. The inputs are generated by randomly sampling numbers from a specific range with some constraints. Noise is introduced by randomly replacing some numbers in the output list with other values. This simulates natural errors, such as data corruption or misrecording, that might occur in sequence-based tasks.
During the data synthesis process, we attach manual supervision to ensure that the generated data can correctly induce the rules. During the evaluation process, we use the inputs in the test set as the input for rule execution. We evaluate the model’s performance using exact match. If the model’s output is correct on all the test set examples, we consider the model to have successfully induced the rule. If the model fails to output a valid programmatic rule or the program contains an infinite loop or errors, we consider it a failure.
@@ B.1 Experimental Settings
For robustness under different noise levels, we run each experiment three times and report the mean and standard deviation of the results to avoid randomness. Except for the self-consistency (SC), hypothesis refinement (HR), and sample-steered rule refinement (SRR) that require diverse generations, we set the temperature to for all models to ensure reproducibility. For SC, HR and SRR, the temperature is set to , consistent with the original work of SC Wang et al. (2023). The positions of noise in seen examples are random to avoid positional bias Lu et al. (2022). In implementation, we choose subsets for SRR by splitting the seen examples into two subsets, and we generate hypotheses at each iteration of HR. The number of iterations is set to for SR, HR and SRR.
@@
For the GPT-4o-mini and GPT-4o models, the experiments are conducted through the official OpenAI333https://openai.com/api/ API platform. For the DeepSeek-V3 and DeepSeek-R1 models, the experiments are conducted through the official DeepSeek444https://platform.deepseek.com/usage API platform and the Volcengine555https://www.volcengine.com/ platform. We spent about USD in total for the experiments.
Method | Acc | Acc | Consistency |
CoT | |||
SC | |||
SR | |||
HR | |||
SRR |
Method Arithmetic7 Arithmetic8 Arithmetic9 Crypto Crypto Crypto List Func. BR BW RW WR BR BW RW WR BR BW RW WR BR BW RW WR BR BW RW WR BR BW RW WR BR BW RW WR GPT-4o-mini DO10% - - - - - - - - - - - - 64 182 22 32 15 258 4 23 0 289 0 11 190 492 45 23 DO20% - - - - - - - - - - - - 51 178 35 36 9 250 10 31 0 296 0 4 160 499 75 16 DO30% - - - - - - - - - - - - 38 188 48 26 14 247 5 34 0 290 0 10 126 493 109 22 GPT-4o DO10% - - - - - - - - - - - - 195 64 10 31 85 137 59 19 7 282 7 4 259 433 40 18 DO20% - - - - - - - - - - - - 180 49 25 46 66 140 78 16 2 283 12 3 220 432 79 19 DO30% - - - - - - - - - - - - 155 56 50 39 49 137 95 19 3 282 11 4 178 433 121 18 CoT 0 94 3 3 2 72 20 6 0 89 3 8 82 13 2 3 7 60 20 13 1 93 3 3 96 131 18 5 SC 0 99 1 0 1 82 12 5 0 97 3 0 83 13 2 2 19 50 21 10 0 88 7 5 100 130 14 6 SR 0 94 5 1 4 64 18 14 0 90 6 4 78 14 5 3 10 56 21 13 2 92 3 3 88 129 22 11 HR 0 94 5 1 12 56 16 16 5 66 15 14 82 14 2 2 40 29 15 16 6 64 14 16 124 91 20 15 SRR 0 89 6 5 23 27 22 28 4 64 17 15 81 14 1 4 39 32 16 13 0 83 8 9 136 96 11 7 DeepSeek-V3 DO10% - - - - - - - - - - - - 65 142 85 8 35 224 30 11 35 213 40 12 282 403 41 24 DO20% - - - - - - - - - - - - 37 141 113 9 37 208 28 27 24 215 51 10 238 400 85 27 DO30% - - - - - - - - - - - - 31 148 119 2 28 213 37 22 19 210 56 15 198 396 125 31 CoT 141 19 26 14 160 2 32 6 110 12 53 25 156 26 13 5 29 114 34 23 5 168 22 5 236 184 56 24 SC 76 7 10 7 92 0 7 1 69 4 15 12 84 11 3 2 25 42 18 15 1 83 10 6 129 91 19 11 SR 58 8 22 12 66 9 17 8 45 15 17 23 64 10 18 8 5 66 15 14 2 83 9 6 105 94 38 13 HR 74 6 12 8 82 2 9 7 67 11 11 11 84 12 2 2 40 28 16 16 4 63 20 13 153 70 21 6 SRR 93 0 4 3 94 0 5 1 88 0 6 6 84 11 3 2 37 32 16 15 2 78 11 9 154 73 15 8
B.2 Additional Results
Different Types of Noise
To ensure the noise reflects perturbations encountered in real-world scenarios, the noise of Arithmetic is different from that of Cryptography and List Functions. To explore the consistency of the results under a more uniform noise strategy, we conducted additional experiments on the base-7 Arithmetic task using DeepSeek-V3. Instead of base-10 outputs, we introduce disturbance noise by adding a random number between and (base_num) to the correct results, aligning this noise more closely with the random corruption used in other tasks. The results are presented in Table 8.
From the results, we find that the performance of CoT, SC, SR and SRR exhibits the same trend as in the original experiments, with a consistent drop in accuracy when noise is introduced. The consistency score also shows a similar trend, indicating the robustness fragility of inductive reasoning in the presence of noise. The consistency drop is more evident than the accuracy drop, suggesting the accuracy drop is not solely due to the failure of previous solvable cases. This further supports the need to consider a more detailed granularity except for the accuracy.
Breakdown of Consistency Score
For a more comprehensive understanding of the consistency score, we provide a detailed breakdown of the consistency score of the experiments in Table 9. From the results, we observe that the introduction of noise not only leads to a performance drop in the original solvable cases but also results in some previously unsolved cases being solved. This reveals the fragility and inconsistency in the rule induction process.
Ablation
We conduct ablation studies to investigate the impact of sampling right and wrong examples during the refinement process of SRR. The results are shown in Table 10. From the results, we observe that sampling right and wrong examples during the refinement process of SRR has a positive impact on the performance when noise is introduced, while it has a slightly negative impact on the performance in the clean condition. This suggests that sampling can help the model to better adapt to the noise, the performance gap between the clean and noisy conditions is reduced.
Dataset Accuracy Accuracy w/sample w/o sample w/sample w/o sample Arithmetic7 Arithmetic8 Arithmetic9 Crypto Crypto Crypto List Func. Average
We also investigate the performance of SR, HR, and SRR across different iterations. The results are shown in Figure 6. Compared to the performance of SRR and HR, the performance of SR is not stable across different refinement iterations, and even shows a performance drop in some cases. This indicates that the LLM-generated feedback in SR is not always helpful for the model to refine the rule induction process. In contrast, both HR and SRR show a more stable performance improvement across iterations, and the improvement is more pronounced in HR. This is because HR generates different refined hypotheses at the same iteration and selects the best one, while SRR only refines the existing hypothesis. Specifically, HR also shows a performance drop in the DeepSeek-V3 model on the List Functions task with noise, which is likely due to HR only selecting the best hypothesis from the generated ones at each iteration. At the same time, SRR chooses the best one from all the generated rules.
Computational Cost
The computational cost of different methods is an important factor to consider the effectiveness, which is measured by the number of output tokens in the whole process. In general, methods that require more samples or generate more responses tend to have higher computational costs. We summarize the computational cost of different methods in Table 11.
Method DeepSeek-V3 DeepSeek-V3 GPT-4o GPT-4o CoT SC SR HR SRR
More Reasoning Models
Except for the reasoning model used in 4.5, we also conduct experiments with QwQ-32B, another reasoning model implementing a reinforcement learning scaling approach Team (2025). Our goal is to explore whether the introduced noise can impact the performance of the reasoning models. The results based on Chain-of-Thought are shown in Table 12. Even the performance of reasoning models is more powerful than the LLMs, they still suffer from the noise and counterfactual tasks.
Dataset DeepSeek-R1 QwQ-32B Acc Acc Consistency Acc Acc Consistency Arithmetic7 Arithmetic8 Arithmetic9 Crypto Crypto Crypto List Func.
Appendix C Prompt Templates
We provide all the prompts we used in this section. For the self-refine, hypothesis refinement, and sample-steered rule refinement setting, we use the chain-of-thought prompt as the initial prompt; the iterative prompts are different due to the type of feedback.